Okay I sort if need help on this word problem, thank you for helping.
(Word problem)
Bob took his quarters and dimes to the bank. He knew he had 500 coins. The teller counted his coins and said his deposit would be for $65.00. How many quarters did Bob deposit.

Answer by mananth(16949)

Quarters---------x numbers
Dimes-----------y numbers

x+y= 500 .............1
25x+10y =6500 .............2 (convert $65 into cents)
Eliminate y
multiply (1)by -10
Multiply (2) by 1
-10x-10y=-5000
25x+10y =6500
Add the two equations
15 x = 1500
/ 15
x = 100
plug value of x in (1)
1 x + 1 y = 500
100 + 1 y = 500
1 y = 500 -100
1 y = 400
y = 400 Quarters 100 numbers
Dimes 400 numbers
m.ananth@hotmail.ca